Rope Man Repair Logs
Repairer: Paul Swan
BOARD NO: 0069
GAME: Rope Man
(1) Dead. No video, no sound.
(2) Continuous thumping noise, other sounds very quiet.
(1) Looking at the pins of the 6809E with a logic probe indicated that all the signals were idle with the reset pin always active. Whilst moving the board set around, the game suddenly started with a sound fault (2). Moving the edge connector reset the game and the 5V rail measured 4.60V at the furthest power extremity, varying depending on how the edge connector was moved. The edge connector itself had burn marks around the two 5V power fingers and the the whole connector layout drifted by about half a finger from one end of the connector to the other. Cleaning the edge fingers with a soft wire brush fixed the game except for (2).
(2) The thumping noise was controlled with the on board volume control suggesting the problem lay behind the amplifier sections. The thumping noise was detectable on the scope as a very low rate ramp on all the audio circuitry right up to the two AY-3-8910's. The thumping noise appeared on all six channels (pins 3,4 & 38 of both devices). Powering up the game with these two devices removed did not eliminate the noise suggesting the problem lay in the filter/pre-amp sections. Using a multimeter to check the pins of the 741 op-amp indicated that the -V input (pin 7) was around +2.5V. This pin was fed via a 1N4002 diode from the -5V supply input at the edge connector and in fact this diode was open circuit. Replacing the diode fixed the sound.